OpenStudy (anonymous):

Write an exponential function y=ab^x for a graph that includes (1, 15) and (0, 6).

OpenStudy (anonymous):

\[y=a\times b^x\] \[6=a\times b^0\] \[6=a\times 1\] \[a = 6\] so we know what a is, now we have to find b

OpenStudy (anonymous):

\[15=6\times b^1\] \[15=6b\] \[b=\frac{15}{6}\] ans we are done

OpenStudy (anonymous):

thank you :)